OpenStudy (anonymous):

A circular grass field has a circumference of 120√π meters. if Eric can mow 400 square meters of grass per hour, how many hours will he take to mow the entire few?

OpenStudy (saifoo.khan):

Steps: Use the circumference given to find the radius. Then use that to find the total area. Then use that area to find the time to be taken by dividing the total area by given area. 3,2,1 Now begin!

OpenStudy (anonymous):

thanks so much!!

OpenStudy (saifoo.khan):

Cheers! (;

OpenStudy (anonymous):

wait but to find the radius do i square the xpression

OpenStudy (saifoo.khan):

The formula is: C = 2 pi*r. or in other words: C = pi *d

OpenStudy (anonymous):

so it has a circumference of 120√π meters so do i square that so remove the root sign or is my diameter 120?

OpenStudy (saifoo.khan):

Nope, stay it that way.

OpenStudy (saifoo.khan):

C = 2* pi*r 120√π = 2* pi * r 60 √π = pi * r Yes you were correct, try it out.

OpenStudy (anonymous):

ohhhh thanks!!!

OpenStudy (saifoo.khan):

Welcome!